# AnchorCrafter: Animate CyberAnchors Saling Your Products via Human-Object Interacting Video Generation

### v1

@@ Abstract
The automatic generation of anchor-style product promotion videos presents promising opportunities in online commerce, advertising, and consumer engagement. However, this remains a challenging task despite significant advancements in pose-guided human video generation. In addressing this challenge, we identify the integration of human-object interactions (HOI) into pose-guided human video generation as a core issue. To this end, we introduce AnchorCrafter, a novel diffusion-based system designed to generate 2D videos featuring a target human and a customized object, achieving high visual fidelity and controllable interactions. Specifically, we propose two key innovations: the HOI-appearance perception, which enhances object appearance recognition from arbitrary multi-view perspectives and disentangles object and human appearance, and the HOI-motion injection, which enables complex human-object interactions by overcoming challenges in object trajectory conditioning and inter-occlusion management. Additionally, we introduce the HOI-region reweighting loss, a training objective that enhances the learning of object details. Extensive experiments demonstrate that our proposed system outperforms existing methods in preserving object appearance and shape awareness, while simultaneously maintaining consistency in human appearance and motion. Project is available at https://cangcz.github.io/Anchor-Crafter/
@@
footnotetext: *Equal contribution.footnotetext: Corresponding author: Fan Tang.
1 Introduction
Viewing unboxing videos and live-streamed product promotions by content creators and live streamers–collectively referred to as anchors–on platforms such as YouTube, TikTok, Douyin, etc. has become an integral part of the online shopping experience for users. Recent advancements in video generation [2, 13, 39, 12, 28] have made it possible to automate the creation of such unboxing and promotional content. However, generating high-fidelity, realistic anchor-style product promotion videos remains a significant technical challenge.
Pose-guided human video generation [32, 14, 42, 36, 4, 41, 16] aligns closely with anchor-style product promotion videos. Existing diffusion-based methods generate temporally consistent, high-fidelity human videos using pose and appearance references. Make-Your-Anchor [16] pioneered personalized anchor-style video generation, while methods like AnimateAnyone [14] and MimicMotion [41] animate static human images with precise poses, adaptable for anchor-style video generation. However, these approaches fail to generate human-driven product displays. As shown in Fig. 2, they treat objects as static textures, preventing appropriate interaction movement.
To address this issue, the main challenge lies in the lack of human-object interaction (HOI) capabilities [37, 5]. Specifically, existing methods fail to understand objects as secondary entities, cannot control object trajectories, and struggle to interpret the complex interactions between human movements and objects. Recent studies on HOI generation [37, 5] primarily focus on hand-centric videos or image domains, which do not offer the degrees of freedom necessary for anchor-style product promotion video generation. Therefore, integrating HOI into pose-guided human video generation remains a critical challenge.
In this study, we propose AnchorCrafter, a novel system for generating anchor-style product promotion videos by integrating human-object interaction (HOI) into pose-guided human video diffusion models. By learning from reference videos, our system automatically creates realistic human-object interaction videos using images of humans and objects, motion sequences, and object trajectories.
AnchorCrafter achieves this through two core components: object appearance preservation and interaction guidance. For appearance preservation, we introduce HOI-appearance perception, which improves object appearance accuracy by fusing multi-view features and using a decoupling network structure to disentangle human-object appearances. For interaction guidance, HOI-motion injection enables precise control of object trajectories using depth and hand 3D mesh input, while mitigating interaction artifacts with occlusion handling strategies. Besides, we use an HOI-region reweighting loss during training to enhance the detail of objects by emphasizing the hand-object interaction region.
(a) Reference Input
(b) AnimateAnyone
(c) MimicMotion
(d) Ours
Figure 2: Animating a reference image of a peron holding an object, existing methods accurately follow human poses but fail to generate hand-object interaction (AnimateAnyone [14]) or misinterpret the object as part of the human, leading to no movement (MimicMotion [41]). Our method accurately generates human-object interactions while preserving the object’s appearance.
We conduct extensive experiments on real-world objects and demonstrate that AnchorCrafter achieves object appearance preserved and interaction awareness results, which are not accomplished by current approaches. Additionally, our method achieves competitive performance with state-of-the-art approaches across various image and video quality evaluations. In summary, our contributions are as follows:
@@ 1 Introduction
  • We propose AnchorCrafter, a novel system incorporating the human-object interaction into pose-guided human video models, and generating realistic anchor-style product promotion videos by animating reference anchors.
@@ 2 Related Work
Pose-Guided Human Video Generation.
@@
Pose-guided human video generation utilized a sequence of poses to control the human motion in the generated videos, commonly applied in scenarios such as dance and speech video generation. With the rapid advancements in diffusion models [11, 26], pose-guided human video human generation archives remarkable progress.
@@
The common approaches leverage ControlNet [40], a parallel control branch attached to the diffusion UNet, to inject human skeletal motion sequences [4, 32, 36, 30, 16], thereby achieving advanced pose guidance. AnimateAnyone [14] proposed ReferenceNet to control the human appearance, which injects the information through the self-attention mechanism. Champ [42] enhances motion control by combining four different types of poses. MimicMotion [41] engineered a lightweight pose guidance network to incorporate human pose conditions, and employed confidence-aware pose guidance to enhance the quality of hand motion synthesis.
@@
These studies demonstrate that by incorporating human poses such as OpenPose [3], DensePose [9], SMPL-X [22], control over human movement can be achieved. However, the important human-object interaction generation is ignored, despite its importance for real-world applications such as anchor-style product promotion video generation, while AnchorCrafter incorporates HOI into human video generation to accomplish it.
Human-Object Interaction Generation.
@@
The goal of Human-Object Interactions Generation (HOI) is to generate visual content that accurately represents the relationships between humans and objects. Some studies primarily focused on HOI motion generation. OMOMO [20], HOI-Diff [24] and InterDiff [35] predict human motions during interactions with objects based on 3D object representations. IMoS [8] generates full-body human and 3D object motions from textual input, but its focus is primarily on hand-grasping small objects.
@@
HOI-Swap [37] zeroes in hand-centric videos, utilizing a two-stage editing technique to swap the object in hand while keeping the hand-object interactions. VirtualModel [5] stands as the most relevant work in literature, albeit within the domain of image generation. It excels in synthesizing interactive images of characters by leveraging input objects and human poses. Existing methods primarily focus on HOI image generation or object swapping in hand-centric videos, leading to the absence of the ability to generate videos with sufficient degrees of freedom, particularly videos that include the entire human body. In contrast, our approach integrates HOI capabilities into a human video generation model, enabling the generation of high-quality, controllable anchor-style product promotion videos.
@@ 3 System Setting
Overview.
Even though existing approaches can generate high-quality human videos from a single image, we argue that for real-world applications like product promotion, a single image as input cannot accurately capture the product’s information for video generation. In this study, we accomplish anchor-style product promotion generation by setting the task as a learning paradigm for customized products: the information of objects is learned from several sample videos, and the model can generate the interaction video with arbitrary human appearance and motion controls.
In practical applications, by capturing interaction guidance provided by an actor and providing only one image of the target anchor, the model can animate the anchor to display the product, which is applicable to crafting real-world unboxing videos or live-streamed product promotions.
Training.
During training, a set of anchor-style product interaction videos is required to depict interactions between humans and specific products. Each source video denoted as $X=\{x^{i:n}\}$ with $n$ frames allows the model to learn the target object $O$ and the interaction distribution between humans and the object. The video duration for each object is configured to exceed twenty seconds to ensure sufficient interaction data. For appearance, a reference human image $I_{H}$ for the target anchor and multi-view images of the target object $I_{O}=\{i_{O}^{1},i_{O}^{2},i_{O}^{3}\}$ are required. We employ multiply comprehensive conditions to control the HOI motions, including the sequences of human skeletons $P=\{p^{1:n}\}$, 3D hand meshes $H=\{h^{1:n}\}$ and depth maps of the object $D=\{d^{1:n}\}$. The system generates a sequence of frames $Y=\{y^{1:n}\}$ with a target human and customized object that captures controlled human-object interactions.
Inference.
In the inference stage, we use HOI motions performed by actors to drive the generation of anchor videos. Actors demonstrate the required interaction actions by holding the products $O$ learned during training, and the extracted control conditions $P$, $H$, $D$ will serve as inputs to the system. Subsequently, our model can drive the generation of product-promotion videos for any unseen anchor image $I_{H}$.
@@ 4 Methodology
The whole pipeline of our proposed system is illustrated in Fig. 3, where two primary components are compromised: the HOI-appearance perception, and the HOI-motion injection. At the beginning of this section, we will briefly introduce the video diffusion model architecture utilized in AnchorCrafter in Sec. 4.1. HOI-appearance perception, which is proposed for appearance control, will be detailed in Sec. 4.2. Next to be introduced in Sec. 4.3 is the HOI-motion injection, where motion guidance with object track and inter-occlusion issues are considered. Furthermore, we adopt an HOI-region reweighting loss aimed at enhancing object details in Sec. 4.4.
@@ 4.1 Video Diffusion Model
AnchorCrafter is based on a video diffusion model [2] architecture, comprising a diffusion UNet [27, 26] with temporal layers, denoted as $\epsilon_{\theta}$, and a variational autoencoder (VAE) [17] composed of an encoder $Enc$ and a decoder $Dec$ for compressing and uncompressing the video frames. During training, the video sequence $X$ is encoded into the latent space as $Z_{0}=Enc(X)=\{z_{0}^{1:n}\}$.
The core training objective for the video diffusion model is:
@@
where $c$ is conditional signal, $Z_{t}=\{z^{1:n}_{t}\}$ is the latent feature diffused from $Z_{0}$ through a deterministic Gaussian process over $t$ timesteps by adding noise $\epsilon$. During inference, the initial noises $U=\{u^{1:n}\}$ sampled from Gaussian noise are denoised with $T$ timesteps to get the estimated $\hat{Z_{0}}$, and the output video with $Y=Dec(\hat{Z_{0}})=\{y^{1:n}\}$.
@@ 4.2 HOI-Appearance Perception
HOI-appearance perception is illustrated in Fig. 4. The design of this module is centered around the extraction of appearance features for both the human and the object, followed by their integration into the backbone network.
We begin with the extraction of human appearance features. We refer to previous approaches [14, 32, 41, 34] by integrating both global context and local details from the input images. Specifically, the VAE encoder $Enc$ maps human images to a latent space compatible with the diffusion model $Z_{H}=Enc(I_{H})$, which preserves the local detail. Then repeat $Z_{H}$ $n$ times and concatenate $Z_{H}$ with the input for each frame $Z_{t}$ of the UNet. Meanwhile, the CLIP [25] features of the human image are extracted as $f_{H}=CLIP(I_{H})$ for global representation.
To accurately depict objects, especially the product being promoted, the model must perceive their shape and texture from multiple perspectives. Instead of a single reference image, we utilize multi-view object reference images $I_{O}=\{i_{O}^{1},i_{O}^{2},i_{O}^{3}\}$ captured from 45° left, front, and 45° right rotations. Similar to human reference, the VAE feature of the front reference $i_{O}^{2}$ is extracted as $Z_{O}=Enc(i_{O}^{2})$ and then repeats and concat with the $Z_{t}$. Different from the process for human reference, we propose multi-view object feature fusion to extract richer information from multi-view images of the input object images.
@@
Multi-View Object Feature Fusion.
We introduce the multi-view object feature fusion to understand object appearance inherently. The process begins by feeding multi-view object reference $I_{O}$ into the pre-trained DINOv2-large model [21]. The three-view DINO features are then processed through two distinct branches, as illustrated in Fig. 4.
To facilitate the model’s learning of the 3D consistency of objects from multi-view features, three-view DINO features are passed through a multi-view self-attention layer. We only retain the output features corresponding to $i_{O}^{2}$, and then another self-attention operation is applied to this tensor. After these self-attention layers, we retained the first twelve dimensions of the features as the final output, with the size of $R^{12\times 1024}$.
In addition, we concatenate the first [CLS] feature from three-view DINO features a vector of size $R^{3\times 1024}$ for global representation, then pass through a linear layer. The two resulting outputs are concatenated as the final multi-view object representation, $f_{O}\in R^{15\times 1024}$. This module is designed to allow the model to exploit the robust representative capability of DINO in extracting 3D consistency features from multi-view object images [29].
Human-Object Dual Adapter.
We observed that directly integrating object features $f_{O}$ into UNet along with human features $f_{H}$ leads to the appearance entanglement of the objects and humans.
We propose a human-object dual adapter by replacing the cross-attention layers in each block of the diffusion UNet to achieve better disentanglement between humans and objects. The human feature $f_{H}$ extracted from CLIP [25] and the object feature $f_{O}$ obtained by multi-view object feature fusion are each fed into a cross-attention layer, and the formula can be written as:

### v2

# AnchorCrafter: Animate Cyber-Anchors Selling Your Products via Human-Object Interacting Video Generation

@@ Abstract
The generation of anchor-style product promotion videos presents promising opportunities in e-commerce, advertising, and consumer engagement. Despite advancements in pose-guided human video generation, creating product promotion videos remains challenging. In addressing this challenge, we identify the integration of human-object interactions (HOI) into pose-guided human video generation as a core issue. To this end, we introduce AnchorCrafter, a novel diffusion-based system designed to generate 2D videos featuring a target human and a customized object, achieving high visual fidelity and controllable interactions. Specifically, we propose two key innovations: the HOI-appearance perception, which enhances object appearance recognition from arbitrary multi-view perspectives and disentangles object and human appearance, and the HOI-motion injection, which enables complex human-object interactions by overcoming challenges in object trajectory conditioning and inter-occlusion management. Extensive experiments show that our system improves object appearance preservation by 7.5% and doubles the object localization accuracy compared to existing state-of-the-art approaches. It also outperforms existing approaches in maintaining human motion consistency and high-quality video generation. Project page including data, code, and Huggingface demo: https://github.com/cangcz/AnchorCrafter.
Index Terms:
Diffusion models, human video generation, digital human synthesis, interaction synthesis.
1 Introduction
Viewing unboxing videos and live-streamed product promotions by content creators and live streamers–collectively referred to as anchors–on platforms such as YouTube, TikTok, Douyin, etc, has become an integral part of the online shopping experience. Recent advancements in computer graphics, including video generation [1, 2] and human animation [3, 4], have made it possible to automate the creation of such unboxing and promotional content. However, achieving high-fidelity temporal consistency, object realism, and controllable motion generation remains a formidable challenge.
Pose-guided human video generation [5, 6, 7, 8, 9, 10] aligns closely with anchor-style product promotion videos. Existing diffusion-based methods generate temporally consistent, high-fidelity human videos using pose and appearance references. Make-Your-Anchor [8] pioneered personalized anchor-style video generation, while methods like MimicMotion [7] and StableAnimator [9] animate static human images with precise poses, adaptable for anchor-style video generation. However, due to the lack of effective human-object interaction (HOI) capabilities [11, 12, 13], these methods are unable to generate human-driven product demonstrations. As illustrated in Fig. 1, objects are often treated as static textures, which restricts interactive realism and hinders the modeling of complex hand-object interactions.
(a) Reference Input
(b) MimicMotion
(c) StableAnimator
(d) Ours
Figure 1: Existing methods accurately follow human poses but struggle with realistic hand-object interactions, often misinterpreting the object as part of the human, leading to static animations. In contrast, our approach ensures natural and dynamic movement by precisely synthesizing human-object interactions while preserving object appearance.
Recent studies on HOI generation [11, 12, 13] primarily focus on hand-centric videos or image domains, which do not offer the degrees of freedom necessary for anchor-style product promotion videos. Therefore, seamlessly integrating HOI into pose-guided human video generation is essential for achieving natural and dynamic human-object interactions in synthesized videos.
@@ 1 Introduction
To generate anchor-style product promotion videos, we introduce AnchorCrafter, a novel object-centric customization framework. Unlike person-centric approaches such as Make-Your-Anchor [8], our method tailors a video diffusion model by collecting a one-minute interaction video of a specific object, enabling arbitrary anchors to naturally interact with that object in diverse poses. Specifically, we refine human and object representations via HOI-appearance perception, which integrates multi-view features and adopts a decoupled architecture to effectively disentangle human and object appearances. However, without precise motion control, objects remain disconnected from human interactions. To achieve natural and coordinated object behavior, we propose the HOI-motion injection module, which leverages depth maps and 3D hand meshes to provide fine-grained motion guidance to the model. In addition, interaction artifacts are alleviated through the adoption of occlusion-handling strategies. Furthermore, conventional training objectives often underrepresent the subtle dynamics of hand-object interactions, which undermines the realism of generated content. To enhance interaction fidelity, we introduce an HOI-region reweighting loss that strategically emphasizes interaction-critical regions during training, leading to reinforced fine-grained object representations and improved realism in modeled interactions. We conduct extensive experiments on real-world objects and demonstrate that AnchorCrafter achieves object appearance preservation and interaction awareness results, which current approaches do not accomplish. Moreover, we collected and open-sourced a digital human interaction dataset to fill the gap in the field of interactive digital human generation. In summary, our contributions are as follows:
@@
  • We propose AnchorCrafter, a novel system incorporating human-object interaction into pose-guided human video models and generating realistic anchor-style product promotion videos by animating reference anchors.
@@ 2 Related Work
2.1 Pose-Guided Human Video Generation
@@
Pose-guided human video generation utilizes sequential poses to control human motion in generated videos, commonly applied in scenarios such as dance and speech video generation. With rapid advancements in diffusion models [14, 15] and pose-controlled human image generation [16, 17], pose-guided human video generation has achieved remarkable progress.
@@
Common practices leverage ControlNet [18], a parallel control branch attached to diffusion UNet, to inject human skeletal motion sequences [19, 6, 20, 8, 3], thereby achieving advanced pose guidance. AnimateAnyone [5] proposed ReferenceNet to control human appearance, which injects information through the self-attention mechanism. Champ [21] enhances motion control by combining four different types of poses. MimicMotion [7] engineered a lightweight pose guidance network to incorporate human pose conditions, and employed confidence-aware pose guidance to enhance the quality of hand motion synthesis. To further improve facial detail embedding, StableAnimator [9] introduces a Face Encoder, enhancing precision in facial synthesis.
Recent studies have explored replacing UNet with diffusion transformer(DiT) to achieve better temporal consistency in video generation. Based on the DiT framework, HumanDiT [22] employs prefix latent reference strategies, ensuring visual consistency across extended sequences. UniAnimate-DiT [23], built upon Wan2.1 [24], incorporates Low-Rank Adaptation [25] to facilitate efficient digital human synthesis while reducing computational overhead. Similarly to UNet-based approaches, both models leverage pose sequences to control human motion within generated images. However, despite these advancements, DiT-based frameworks still require large-scale datasets for training to uphold video quality and coherence, posing challenges in practical deployment.
@@
These studies demonstrate that by incorporating human poses such as OpenPose [26], DensePose [27], SMPL-X [28], control over human movement can be achieved. However, the important human-object interaction generation is often overlooked, despite its vital importance for real-world applications such as anchor-style product promotion video generation. In contrast, AnchorCrafter incorporates HOI into human video generation to address this issue.
2.2 Human-Object Interaction Generation
@@
The goal of HOI is to generate visual content that accurately represents the relationships between humans and objects. In the domain of 3D reconstruction, EasyHOI [29] enables hand-object interaction reconstruction from a single image, whereas DICE [30] facilitates single-image reconstruction of hand-face interactions. In the field of motion generation [31, 32], some studies focus on interactive motion synthesis, aiming to generate human-object interactions with enhanced realism. InterDiff [33], TeSMo [34] and HOIAnimator [35] predict human motions during interactions with objects based on 3D object representations. PMP [36] learns to use part-wise motion priors to physically interact with environments. IMoS [37] generates full-body human and 3D object motions from textual input, but its focus is primarily on hand-grasping small objects. TextIM synthesizes human interactive motions with part-level semantic accuracy by aligning textual descriptions with interactive body movements.
@@
However, human-object interaction remains relatively underexplored in the field of content generation. HOI-Swap [11] zeroes in hand-centric videos, utilizing a two-stage editing technique to swap the object in hand while keeping the hand-object interactions. ReHold [13]proposes the HOI Restoration Module, which injects hand and object details into the model to enhance hand-object interaction. VirtualModel [12] stands as the most relevant work in literature, albeit within the domain of image generation. It excels in synthesizing interactive images of characters by leveraging input objects and human poses. Existing methods primarily focus on HOI image generation or object swapping in hand-centric videos, leading to the absence of the ability to generate videos with sufficient degrees of freedom, particularly videos that include the entire human body. In contrast, our approach integrates HOI capabilities into a human video generation model, enabling the generation of high-quality, controllable anchor-style product promotion videos.
@@ 3 System Setting
Figure 3: Fine-tuning for new products and inference for new anchors. Fine-tuning the model with a one-minute video to achieve a customized model for new objects. After fine-tuning, our method could generate arbitrary anchor videos selling the product with various unseen motions.
Overview. Driven by cutting-edge progress in pose estimation [38, 39] and enhanced by access to large-scale human datasets [40, 41, 42, 43], techniques [7, 19, 6, 41] for producing high-fidelity, controllable human videos have become increasingly mature. However, the absence of open-source human-object interaction datasets poses a significant challenge to generating anchor-style product promotion videos requiring accurate reconstruction of objects. To better address customized product interactions, we frame the task as a learning paradigm for customized products: learns fine-grained texture and appearance from a one-minute interaction video with the product. By leveraging this learned representation, the model generalizes effectively, producing interaction videos with diverse human appearances and precisely controlled motions. The learning framework of our system consists of two stages: training on diverse datasets to capture the distribution of HOI videos, and fine-tuning on specific objects, enabling more precise and adaptable interaction modeling.
Training. During training, each source video $X=\{x^{1:f}\}$ with $f$ frames enables the model to learn the target object $O$ and its interaction patterns with humans. To ensure accurate appearance modeling, we use a reference image $I_{H}$ for the anchor and multi-view images $I_{O}=\{i_{O}^{1},i_{O}^{2},i_{O}^{3}\}$ for the object. We employ multiply comprehensive conditions to control the HOI motions, including the sequences of human skeletons $P=\{p^{1:f}\}$, 3D hand meshes $H=\{h^{1:f}\}$ and depth maps of the object $D=\{d^{1:f}\}$. The system then generates a sequence of frames $Y=\{y^{1:f}\}$ featuring a target human and customized object with controlled interactions. To improve interaction modeling, we collect a foundational dataset of human-object interactions. Given the diversity of objects and dataset constraints, our primary focus is refining interaction control rather than capturing intricate texture details.
Fine-tuning. During fine-tuning, we collect videos exceeding one minute in length for each object requiring demonstration, as illustrated in Fig. 3. Benefiting from the training phase, where the model learns human-object interaction concepts, the fine-tuning stage requires only a short video to grasp object textures. This enables the model to generate previously unseen poses and novel anchor interactions. We maintain that collecting one-minute videos is both practical and contributes to a more realistic representation of the object.
Inference. In the inference stage, we use HOI motions performed by actors to drive the generation of anchor videos. Actors demonstrate the required interaction actions by holding the products $O$ learned during fine-tuning, and the extracted control conditions $P$, $H$, and $D$ will serve as inputs to the system. Subsequently, our model can generate product-promotion videos for any unseen anchor image $I_{H}$.
@@ 4 Methodology
As illustrated in Fig. 4, our system’s pipeline consists of two core components: HOI-appearance perception and HOI-motion injection. We first introduce the video diffusion model architecture utilized in AnchorCrafter in Sec. 4.1. To enhance the reconstruction of both human and object appearances, we design HOI-appearance perception (Sec. 4.2), which employs an appearance-decoupling strategy to ensure accurate representation. To regulate the movement of humans and objects, we integrate HOI-motion injection (Sec. 4.3), which introduces precise control signals into the system. Additionally, to refine interaction fidelity, we introduce HOI-region reweighting loss (Sec. 4.4), which strategically enhances hand-object interaction details during inference.
@@ 4.1 Video Diffusion Model
AnchorCrafter is based on a video diffusion model [44] architecture, comprising a diffusion UNet [45, 15] with temporal layers, denoted as $\epsilon_{\theta}$, and a variational autoencoder (VAE) [46] composed of an encoder $Enc$ and a decoder $Dec$ for compressing and uncompressing the video frames. During training, the video sequence $X$ is encoded into the latent space as $Z_{0}=Enc(X)=\{z_{0}^{1:f}\}$. The core training objective for the video diffusion model is:
@@
where $c$ is conditional signal, $Z_{t}=\{z^{1:f}_{t}\}$ is the latent feature diffused from $Z_{0}$ through a deterministic Gaussian process over $t$ timesteps by adding noise $\epsilon$. During inference, the initial noises $U=\{u^{1:f}\}$ sampled from Gaussian noise are denoised with $T$ timesteps to get the estimated $\hat{Z_{0}}$, and the output video with $Y=Dec(\hat{Z_{0}})=\{y^{1:f}\}$.
@@ 4.2 HOI-Appearance Perception
As illustrated in Fig. 5, the HOI-appearance perception design focuses on extracting appearance features for humans and objects, which are then integrated into the backbone network. Referring to previous approaches [5, 19, 7, 47], we integrate global context and local details from input images for human appearance.
Specifically, the VAE encoder $Enc$ maps human images to a latent space compatible with the diffusion model $Z_{H}=Enc(I_{H})$, which preserves the local detail. Subsequently, $Z_{H}$ is repeated $f$ times and concatenated with the UNet input $Z_{t}$. Meanwhile, the CLIP [48] features of the human image are extracted as $f_{H}=CLIP(I_{H})$ for global representation.
To accurately depict objects, especially the promoting product, the model must perceive their shape and texture from multiple perspectives. Instead of a single reference image, we utilize multi-view object reference images $I_{O}=\{i_{O}^{1},i_{O}^{2},i_{O}^{3}\}$ captured from 45° left, front, and 45° right angles. Similar to human reference, the VAE feature of the front reference $i_{O}^{2}$ is extracted as $Z_{O}=Enc(i_{O}^{2})$ and then repeated $f$ times and concatenated with $Z_{t}$. Different from the process for human reference, we propose multi-view object feature fusion to extract richer information from multi-view images of the input object images.
@@
4.2.1 Multi-View Object Feature Fusion
We introduce the multi-view object feature fusion to understand object appearance inherently. The process starts by feeding the multi-view object reference $I_{O}$ into the pre-trained DINOv2-large model [49], resulting in the extraction of embedding $E_{O}=\{e_{O}^{1},e_{O}^{2},e_{O}^{3}\}\in R^{n\times m},n=1370,m=1024$. Then, $E_{O}$ is processed through two distinct branches, as illustrated in Fig. 5. In the first branch, we designed a multi-view self-attention layer to merge object features from three perspectives, where three object embeddings of shape $n\times m$ from DINO are concatenated along the $n$ channel and then a self-attention is performed. Subsequently, we extract the $e_{O}^{2}$ of the embedding, perform self-attention, and employ a linear network to project the features into a reduced dimensional space of $16\times m$. In the other branch, we concatenate the first [CLS] feature from $E_{O}$ as a vector of size $R^{3\times m}$ for global representation, then pass through a linear layer. Finally, concatenate two branches’ outputs as the final multi-view object representation, $f_{O}\in R^{19\times m}$. This module extracts object features precisely from images captured from three perspectives, extracts 3D consistency features from multi-view images [50], and facilitates the final reproduction of objects.
4.2.2 Human-Object Dual Adapter
We observed that directly integrating object features $f_{O}$ into UNet along with human features $f_{H}$ leads to the appearance entanglement of the objects and humans. We propose a human-object dual adapter by replacing the cross-attention layers in each block of the diffusion UNet to achieve better disentanglement between humans and objects. The human feature $f_{H}$ extracted from CLIP [48] and the object feature $f_{O}$ obtained by multi-view object feature fusion are each fed into a cross-attention layer, and the formula can be written as: